A toy gun uses a spring with a force constant of 290 N/m to propel a 8.5-g steel ball. Assuming the spring is compressed 6.8 cm and friction is negligible, answer the following questions:

What is the gun’s maximum range, in meters, on level ground?

Why does the gun's maximum gun range = (v)^2/g?
Answer #1

Тех. 1-н kxa H = 2 2 H= 290 x 10.0687² 24 0.0085 * 9.7 Ha es 8.30 m Answer = 8.30 m - Ikx² = 1 m² so v = 4 + 295 i ns - . So